28 July 2008 LOIRP Status

Dennis Wingo: The project is rolling on. We have had some good progress on many fronts. Today we had a telecon with JPL (Rafael Alanis and Steven Adams) about the metadata tagging for the images produced from this effort. It was a good telecon and we expect to receive the templates and other information that will allow us to put together the images in the format that will properly integrate with the Planetary Data System.
We also had a major breakthrough today from a friend of a friend of a friends friend for finding documentation for the Ampex FR-900 that is our best drive that we have been working on and operating for the last couple of weeks. A former field service manager from Ampex was found who has the critical schematics and assembly drawings stored on Aperture cards (microfilm on a Hollerith punch card). We have sent him our assembly numbers from our equipment and he is getting a quotation for printing and digitizing the data. We will also have the one document from the Stanford Ampex archive by early next week. Therefore we are reasonably confident that we will have sufficient documentation to fully return to operation the primary FR-900A drive that we have been working on.
Today though, since Ken is troubleshooting individual circuits on the drive and this is not amenable to three or four people working on that simultaneously, I have tasked the students to shift around working with Ken and to completely disassemble and clean the second best drive, an Ampex 901 (that we now know was designed and built especially for the USAF and Eglin AFB) drive that we already have all the documentation on. We have repaired this drive with the top frame from another drive (pictures enclosed) and have washed the drives power supplies and other subsystems. We will attempt to return it to minimal operational status this week as we have time, while focusing on the primary FR-900A drive.
If the follow on funding is obtained to refurbish the heads and the operations, it is our hope that we might possibly have two operational drives (NO PROMISES) which would shorten the time to digitize the images. We will know within a week or two whether this path makes any sense.
Troubleshooting is beginning in ernest now with Ken tracing out circuits. He has found a couple of problems that have been fixed and more that are in process. This is just a very detailed, circuit by circuit process. We have ordered several thousand dollars worth of parts in the last few days and some have already been received and installed. Note in the latest pictures the blue capacitors that are replacing the silver ones. The newer ones are generally much smaller for the same capacitance so some machining of sleeves to make up the difference will be required. Pictures follow.
Tomorrow is tape inventory day with Keith Cowing leading the inventory team of students to get an accurate count of the tapes and then to sort and properly label all the tapes so that we can find the tape that we want when we want during the digitizing process.
All in all we are making steady progress and have reached all of our month one milestones and are well on the way with month two work.